how many sides would a regular polygon have if each side was 170 degrees
Answer by Alan3354(21583)

how many sides would a regular polygon have if each side was 170 degrees
--------------
I don't think that can be answered.
-----------------
However, if each angle is 170 degs:
The total of angles in a polygon of n sides is 180*(n-2)
Each angle is 180*(n-2)/n
180*(n-2)/n = 170
180n - 360 = 170n
-360 = -10n
n = 36 sides (and 36 angles)